one month revenue from the 95 repairs at an auto shop totaled $10,020. The average charges for the repairs on personal and commercial vehicles were $125 and $90 respectively. Find the number of each vehicle repaired.
Answer by stanbon(75887)

one month revenue from the 95 repairs at an auto shop totaled $10,020. The average charges for the repairs on personal and commercial vehicles were $125 and $90 respectively. Find the number of each vehicle repaired.
----
Quantity::: A + B = 95 repairs
Value::::: 90A +125B = 10,020 dollars
-----
Multiply thru the top equation by 90
90A + 90B = 90*95
90A +125B = 10,020
----
Subtract and solve for "B"
35B = 1470
B = 42 (# of $125 repairs)
----
Solve for "A:
A + B = 95
A + 42 = 95
A = 53 (# of $90 repairs)
=============================
Cheers,
Stan H.
================